OpenStudy (anonymous):

4. What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250

OpenStudy (amistre64):

1-r^n ----- but that would imply we knew r, or at least would have to find r 1- r

OpenStudy (amistre64):

should we divide out the 10 first?

OpenStudy (amistre64):

78125 is left and that might indicate a base 5?

OpenStudy (anonymous):

divide 8th term by first gives r^7 so can find r this way

OpenStudy (amistre64):

5^7 = 78125 :)

OpenStudy (anonymous):

I'm working on the problem now. It might take me a minute or two.

OpenStudy (anonymous):

r = 5

OpenStudy (amistre64):

\[A = a +ar^2+ar^3+...+ar^7\] \[a*[A=1+r+r^2+r^3+...+r^7]\]

OpenStudy (anonymous):

Yes, I also got r = 5.

OpenStudy (anonymous):

i make it 976560

OpenStudy (amistre64):

\[10*\frac{1-5^8}{1-5}=sum\] perhaps?

OpenStudy (anonymous):

r=5 s=10(1-5to 8)/4=976560

OpenStudy (anonymous):

Yes, sorry I'm so slow. This is the same answer I got :)